# Yirui Zhou, Xiaowei Liu, Xiaofeng Zhang, Yangchun Zhang

### v1

@@ 7 Conclusion and discussion
In addition to adversarial training, OT is also commonly employed for reward calculation (Arjovsky et al.,, 2017; Papagiannis and Li,, 2022; Luo et al.,, 2023; Liu et al.,, 2024). By integrating distributional RL with OT methods, modeling the uncertainty in the reward distribution has the potential to enhance the robustness of solutions, especially in environments characterized by sparse or noisy reward signals. This is an area we intend to explore in the future.
@@ A.1 Proof of Theorem 1
We first focus on the boundedness of $\sup_{r\in\mathcal{R}}\left|\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi_{\rm E}}}{[r(s,s^{\prime})]}-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mu}_{\pi_{\rm E}}}{[r(s,s^{\prime})]}\right|$. Employing McDiarmid’s inequality (Mohri et al.,, 2018), then with probability at least $41-\delta/$ 4, we obtain that
@@
Connecting the Rademacher complexity to its empirical version, then with probability at least $41-\delta/$ 4, we get that
@@
^$\hat{\mathfrak{R}}_{\mu_{\pi_{\rm E}}}^{(n)}(\mathcal{R})=\mathbb{E}_{\boldsymbol{\sigma}}\left[\sup_{r\in\mathcal{R}}\sum_{i=1}^{n}\frac{1}{n}\sigma_{i}r(s^{(i)},s^{\prime(i)})\right].$
Combining Eq. (A.2) with Eqs. (A.3) and (A.4), with probability at least $21-\delta/$ 2, we obtain that
@@ A.2 Proof of Theorem 2
$~{}~{}~{}~{}\bm{e}(C_{r_{\rm I}}\mu_{\pi_{\rm E}},C_{r_{\rm I}}\mu_{\Pi})-\bm{e}(\hat{c}_{r_{\rm I}}^{(m)}\mu_{\pi_{\rm E}},\hat{c}_{r_{\rm I}}^{(m)}\mu_{\Pi})$
$=C_{r_{\rm I}}\inf_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big{[}\mu_{\pi_{\rm E}}(s,s^{\prime})-\mu_{\pi}(s,s^{\prime})\big{]}\right\}-\hat{c}_{r_{\rm I}}^{(m)}\inf_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big{[}\mu_{\pi_{\rm E}}(s,s^{\prime})-\mu_{\pi}(s,s^{\prime})\big{]}\right\}$
$\overset{(i)}{\geq}C_{r_{\rm I}}\inf_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big{[}\mu_{\pi_{\rm E}}(s,s^{\prime})-\mu_{\pi}(s,s^{\prime})\big{]}\right\}-C_{r_{\rm I}}\inf_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big{[}\mu_{\pi_{\rm E}}(s,s^{\prime})-\mu_{\pi}(s,s^{\prime})\big{]}\right\}$
$\geq C_{r_{\rm I}}\inf_{\pi\in\Pi}\Big{\{}\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big{[}\mu_{\pi_{\rm E}}(s,s^{\prime})-\mu_{\pi}(s,s^{\prime})\big{]}-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big{[}\mu_{\pi_{\rm E}}(s,s^{\prime})-\mu_{\pi}(s,s^{\prime})\big{]}\Big{\}}$
$\geq C_{r_{\rm I}}\inf_{\pi\in\Pi}\Big{\{}\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big{[}\mu_{\pi_{\rm E}}(s,s^{\prime})\big{]}-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big{[}\mu_{\pi_{\rm E}}(s,s^{\prime})\big{]}\Big{\}}+C_{r_{\rm I}}\inf_{\pi\in\Pi}\Big{\{}\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big{[}\mu_{\pi}(s,s^{\prime})\big{]}-\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big{[}\mu_{\pi}(s,s^{\prime})\big{]}\Big{\}}$
$=\Big{(}\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi_{\rm E}}(s,s^{\prime})\big{]}-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi_{\rm E}}(s,s^{\prime})\big{]}\Big{)}-\sup_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big{]}-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big{]}\right\},$ (A.7)
@@
We first focus on the boundedness of $\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi_{\rm E}}(s,s^{\prime})\big{]}-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi_{\rm E}}(s,s^{\prime})\big{]}$. Employing McDiarmid’s inequality, then with probability at least $31-\delta/$ 3, we have that
@@
We then investigate the boundedness of $\sup_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big{]}-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big{]}\right\}$. By McDiarmid’s inequality, then with probability at least $31-\delta/$ 3, we obtain that
$~{}~{}~{}~{}\sup_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big{]}-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big{]}\right\}$
$\leq\mathbb{E}\left[\sup_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big{]}-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big{]}\right\}\right]+2B_{\Pi}\sqrt{\frac{\log(3/\delta)}{2m}},$ (A.9)
where the external expectation is over the random sampling of the distribution ^$\hat{\mathbb{R}}_{\rm I}$ with $m$ state-transition pairs. Based on the Rademacher complexity theory, the first term in Eq. (A.9) can be bounded by
$\hskip-25.6073pt~{}~{}~{}~{}\mathbb{E}\left[\sup_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big{]}-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big{]}\right\}\right]$
$\hskip-25.6073pt\leq 2\mathbb{E}_{\boldsymbol{\sigma},\mathbb{R}_{\rm I}}\left[\sup_{\pi\in\Pi}\sum_{j=1}^{m}\frac{1}{m}\sigma_{j}C_{r_{\rm I}}\mu_{\pi}(s^{(j)},s^{\prime(j)})\right]$
$\hskip-25.6073pt=2\mathfrak{R}_{\mathbb{R}_{\rm I}}^{(m)}(C_{r_{\rm I}}\mu_{\Pi}).$ (A.10)
Connecting the Rademacher complexity to its empirical version, then with probability at least $31-\delta/$ 3, we get that
@@
^$\hat{\mathfrak{R}}_{\mathbb{R}_{\rm I}}^{(m)}(C_{r_{\rm I}}\mu_{\Pi})=\mathbb{E}_{\boldsymbol{\sigma}}\left[\sup_{\pi\in\Pi}\sum_{j=1}^{m}\frac{1}{m}\sigma_{j}C_{r_{\rm I}}\mu_{\pi}(s^{(j)},s^{\prime(j)})\right].$
Combining Eq. (A.9) with Eqs. (A.10) and (A.11), with probability at least $31-2\delta/$ 3, we obtain that
$\sup_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big{]}-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big{[}C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big{]}\right\}\leq 2\hat{\mathfrak{R}}_{\mathbb{R}_{\rm I}}^{(m)}(C_{r_{\rm I}}\mu_{\Pi})+6B_{\Pi}\sqrt{\frac{\log(3/\delta)}{2m}}.$ (A.12)

### v2

@@ 7 Conclusion and discussion
In addition to adversarial training, OT is also employed for reward calculation (Arjovsky et al.,, 2017; Papagiannis and Li,, 2022; Luo et al.,, 2023; Liu et al.,, 2024). By integrating distributional RL with OT methods, modeling the uncertainty in the reward distribution has the potential to enhance the stability of the training process, especially in environments characterized by sparse or noisy reward signals. We regard this area to be explored in the future.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 12301351.
Appendix
@@ A.1 Proof of Theorem 1
We first focus on the boundedness of
$\sup_{r\in\mathcal{R}}\left|\mathbb{E}_{(s,s^{\prime})\sim\mu_{\pi_{\rm E}}}{[r(s,s^{\prime})]}-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mu}_{\pi_{\rm E}}}{[r(s,s^{\prime})]}\right|.$
Employing McDiarmid’s inequality (Mohri et al.,, 2018), then with probability at least $1-\delta/4$, we obtain that
@@
Connecting the Rademacher complexity to its empirical version, then with probability at least $1-\delta/4$, we get that
@@
$\hat{\mathfrak{R}}_{\mu_{\pi_{\rm E}}}^{(n)}(\mathcal{R})=\mathbb{E}_{\boldsymbol{\sigma}}\left[\sup_{r\in\mathcal{R}}\sum_{i=1}^{n}\frac{1}{n}\sigma_{i}r(s^{(i)},s^{\prime(i)})\right].$
Combining Eq. (A.2) with Eqs. (A.3) and (A.4), with probability at least $1-\delta/2$, we obtain that
@@ A.2 Proof of Theorem 2
$\thinspace~~~~\bm{e}(C_{r_{\rm I}}\mu_{\pi_{\rm E}},C_{r_{\rm I}}\mu_{\Pi})-\bm{e}(\hat{c}_{r_{\rm I}}^{(m)}\mu_{\pi_{\rm E}},\hat{c}_{r_{\rm I}}^{(m)}\mu_{\Pi})$
$\thinspace=C_{r_{\rm I}}\inf_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big[\mu_{\pi_{\rm E}}(s,s^{\prime})-\mu_{\pi}(s,s^{\prime})\big]\right\}-\hat{c}_{r_{\rm I}}^{(m)}\inf_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big[\mu_{\pi_{\rm E}}(s,s^{\prime})-\mu_{\pi}(s,s^{\prime})\big]\right\}$
$\thinspace\overset{(i)}{\geq}C_{r_{\rm I}}\inf_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big[\mu_{\pi_{\rm E}}(s,s^{\prime})-\mu_{\pi}(s,s^{\prime})\big]\right\}-C_{r_{\rm I}}\inf_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big[\mu_{\pi_{\rm E}}(s,s^{\prime})-\mu_{\pi}(s,s^{\prime})\big]\right\}$
$\thinspace\geq C_{r_{\rm I}}\inf_{\pi\in\Pi}\Big\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big[\mu_{\pi_{\rm E}}(s,s^{\prime})-\mu_{\pi}(s,s^{\prime})\big]-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big[\mu_{\pi_{\rm E}}(s,s^{\prime})-\mu_{\pi}(s,s^{\prime})\big]\Big\}$
$\thinspace\geq C_{r_{\rm I}}\inf_{\pi\in\Pi}\Big\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big[\mu_{\pi_{\rm E}}(s,s^{\prime})\big]-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big[\mu_{\pi_{\rm E}}(s,s^{\prime})\big]\Big\}+C_{r_{\rm I}}\inf_{\pi\in\Pi}\Big\{\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big[\mu_{\pi}(s,s^{\prime})\big]-\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big[\mu_{\pi}(s,s^{\prime})\big]\Big\}$
$\thinspace=\Big(\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi_{\rm E}}(s,s^{\prime})\big]-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi_{\rm E}}(s,s^{\prime})\big]\Big)-\sup_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big]-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big]\right\},$ (A.7)
@@
We first focus on the boundedness of
$\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi_{\rm E}}(s,s^{\prime})\big]-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi_{\rm E}}(s,s^{\prime})\big].$
Employing McDiarmid’s inequality, then with probability at least $1-\delta/3$, we have that
@@
We then investigate the boundedness of
$\sup_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big]-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big]\right\}.$
By McDiarmid’s inequality, then with probability at least $1-\delta/3$, we obtain that
$\thinspace~~~~\sup_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big]-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big]\right\}$
$\thinspace\leq\mathbb{E}\left[\sup_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big]-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big]\right\}\right]+2B_{\Pi}\sqrt{\frac{\log(3/\delta)}{2m}},$ (A.9)
where the external expectation is over the random sampling of the distribution $\hat{\mathbb{R}}_{\rm I}$ with $m$ state-transition pairs. Based on the Rademacher complexity theory, the first term in Eq. (A.9) can be bounded by
$\thinspace~~~~\mathbb{E}\left[\sup_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big]-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big]\right\}\right]$
$\thinspace\leq 2\mathbb{E}_{\boldsymbol{\sigma},\mathbb{R}_{\rm I}}\left[\sup_{\pi\in\Pi}\sum_{j=1}^{m}\frac{1}{m}\sigma_{j}C_{r_{\rm I}}\mu_{\pi}(s^{(j)},s^{\prime(j)})\right]$
$\thinspace=2\mathfrak{R}_{\mathbb{R}_{\rm I}}^{(m)}(C_{r_{\rm I}}\mu_{\Pi}).$ (A.10)
Connecting the Rademacher complexity to its empirical version, then with probability at least $1-\delta/3$, we get that
@@
$\hat{\mathfrak{R}}_{\mathbb{R}_{\rm I}}^{(m)}(C_{r_{\rm I}}\mu_{\Pi})=\mathbb{E}_{\boldsymbol{\sigma}}\left[\sup_{\pi\in\Pi}\sum_{j=1}^{m}\frac{1}{m}\sigma_{j}C_{r_{\rm I}}\mu_{\pi}(s^{(j)},s^{\prime(j)})\right].$
Combining Eq. (A.9) with Eqs. (A.10) and (A.11), with probability at least $1-2\delta/3$, we obtain that
$\thinspace~~~~\sup_{\pi\in\Pi}\left\{\mathbb{E}_{(s,s^{\prime})\sim\mathbb{R}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big]-\mathbb{E}_{(s,s^{\prime})\sim\hat{\mathbb{R}}_{\rm I}}\big[C_{r_{\rm I}}\mu_{\pi}(s,s^{\prime})\big]\right\}\leq 2\hat{\mathfrak{R}}_{\mathbb{R}_{\rm I}}^{(m)}(C_{r_{\rm I}}\mu_{\Pi})+6B_{\Pi}\sqrt{\frac{\log(3/\delta)}{2m}}.$ (A.12)